### v1

# Leveraging LLMs for Enhanced Open-Vocabulary 3D Scene Understanding in Autonomous Driving

Abstract
This paper introduces a novel method for open-vocabulary 3D scene understanding in autonomous driving by combining Language Embedded 3D Gaussians with Large Language Models (LLMs) for enhanced inference. We propose utilizing LLMs to generate contextually relevant canonical phrases for segmentation and scene interpretation. Our method leverages the contextual and semantic capabilities of LLMs to produce a set of canonical phrases, which are then compared with the language features embedded in the 3D Gaussians. This LLM-guided approach significantly improves zero-shot scene understanding and detection of objects of interest, even in the most challenging or unfamiliar environments. Experimental results on the WayveScenes101 dataset demonstrate that our approach surpasses state-of-the-art methods in terms of accuracy and flexibility for open-vocabulary object detection and segmentation. This work represents a significant advancement towards more intelligent, context-aware autonomous driving systems, effectively bridging 3D scene representation with high-level semantic understanding.
keywords:
Semantic Scene Understanding, Language-Embedded3D Gaussians, LLM, Autonomous Driving
Figure 1: Overview of our LLM-enhanced Language Embedded 3D Gaussian Splatting pipeline. The LE3DGS model generates language feature maps, which are then processed using Algorithm 1 to compute relevancy scores with LLM-generated queries ($p_{\texttt{quer}}$), helping positives ($p_{\texttt{help}}$), and canonical phrases ($p_{\texttt{canon}}$). The system responds to high-level driving scenario queries by highlighting relevant objects (cars, traffic lights, pedestrians) across multiple views.
1 Introduction
Understanding complex scenes in three dimensions is crucial for advancements in autonomous systems, augmented reality, and robotics. The field of 3D scene understanding has seen significant advancements in recent years, particularly with the emergence of neural radiance fields (NeRF) [1] and 3D Gaussian Splatting (3DGS) [2] for novel view synthesis. However, the challenge of open-vocabulary querying in 3D space remains critical for tasks such as object localization and segmentation in autonomous driving scenarios.
Concurrently, large language models (LLMs) [3, 4, 5, 6, 7] have demonstrated remarkable abilities in grasping context and effectively analyzing and interpreting complex scenarios [8]. Recent works have begun to explore the integration of LLMs into autonomous driving tasks. For instance, NuPrompt [9], a large-scale dataset for 3D object detection and tracking in autonomous driving, utilizes natural language prompts to predict object trajectories across different views and frames. While this approach outperforms traditional methods, it primarily uses LLMs to enhance 3D detection tasks rather than directly influencing tracking [8]. Further advancements in this direction include HiLM-D [10], which leverages multimodal LLMs (MLLMs) [11, 12, 13] to improve autonomous driving tasks by integrating high-resolution information. Similarly, ToKEN [14] introduces a novel MM-LLM that tokenizes the world into object-level knowledge, enabling better utilization of LLM’s reasoning capabilities to enhance autonomous vehicle planning in long-tail scenarios.
Building upon these developments, our study proposes a novel method that combines LLMs with Language Embedded 3D Gaussians (LE3DGS) to improve autonomous vehicle scene understanding. This approach addresses the limitations of existing methods in open-vocabulary segmentation and retrieval. The key innovation lies in using LLMs to enhance text embeddings at inference with contextual understanding, enabling a more sophisticated and flexible perception system capable of interpreting complex environmental cues. Figure 1 illustrates the pipeline of our proposed approach.
By integrating the contextual reasoning capabilities of LLMs with the spatial precision of 3D Gaussians, our method aims to significantly improve the understanding of driving scenes. This integration not only enhances the system’s ability to identify and localize objects but also provides a richer semantic interpretation of the scene. The main contributions of our work are as follows:
  • We propose a novel integration of LLMs with LE3DGS for enhanced scene understanding in autonomous driving.
  • We demonstrate how LLMs can be used to generate contextually relevant queries and canonical phrases, improving the flexibility and adaptability of 3D scene analysis.
  • We present experimental results on the Wayvescene101 dataset, showcasing improved performance in both accuracy and adaptability of object detection and scene understanding compared to existing methods.
2 Related Work
Recent work has focused on incorporating semantic information into 3D scene representations. In [15], authors extended neural radiance fields (NeRF) to jointly encode semantics with appearance and geometry, enabling 2D semantic labels using a small amount of in-place annotations. Neural Feature Fusion Fields (N3F) [16] distills 2D image features into a 3D student network, improving dense 2D image feature extractors when applied to multiple images reconstructible as a 3D scene. Language Embedded Radiance Fields (LERF) [17] grounds language embeddings from models like CLIP [18] into NeRF, enabling open-ended language queries in 3D. This approach allows for pixel-aligned, zero-shot queries on distilled 3D CLIP embeddings without relying on region proposals or masks.
3D Gaussian Splatting [2] has emerged as an efficient method for high-quality novel view synthesis. It offers a rendering quality comparable to NeRF but with faster training and inference times [19]. Feature 3DGS [20] enables 3D Gaussian splatting on arbitrary-dimension semantic features via 2D foundation model distillation. This method showcases novel view semantic segmentation, language-guided editing, and segmenting anything through learning feature fields from state-of-the-art 2D foundation models. LE3DGS [21] addresses the challenge of embedding language features into 3D Gaussians without prohibitive memory usage. This approach uses a dedicated quantization scheme and a novel embedding procedure to achieve smooth yet accurate queries while countering multi-view feature inconsistencies. Authors in [22] presented ConceptFusion, an open-set multimodal 3D mapping approach that leverages foundation models pre-trained on internet-scale data to reason about concepts across modalities such as natural language, images, and audio. This method enables effective zero-shot spatial reasoning without additional training or fine-tuning. FeatureNeRF [23] learns generalizable NeRFs by distilling pre-trained vision foundation models. This approach allows mapping 2D images to continuous 3D semantic feature volumes, which can be used for various downstream tasks such as semantic key point transfer and object part segmentation.
Despite these advancements, methods such as LE3DGS [21], LERF [17], and LangSplat [24] rely on a fixed set of predefined canonical phrases during inference. This constraint limits their adaptability to novel or context-specific scenarios, particularly in the dynamic and unpredictable environment of autonomous driving. Moreover, CLIP’s text input length is significantly constrained due to its text encoder being limited to 77 tokens, primarily trained on short texts. This results in a reduced effective token length which has been shown that the actual usable length is just 20 tokens [25]. This limits the previous method’s inference capabilities to very short queries.
Our work addresses this limitation by introducing a novel approach that leverages LLMs to generate queries and canonical phrases at inference time. By doing so, we enable a more flexible and context-aware system that can adapt its language understanding to the specific nuances of each scene. This approach allows for a more accurate and comprehensive interpretation of the 3D environment, capturing subtle details and rare scenarios that might be overlooked by systems relying on predefined canonical language sets. Through this integration of LLMs with LE3DGS, we aim to significantly enhance the open-vocabulary capabilities and overall scene understanding in autonomous driving contexts.
3 Methodology
3.1 Scene Representation with 3D Gaussians
We use the LE3DGS approach [21] for generating 3D Gaussians with language features integrated. This method represents the scene using language embedded Gaussian points and allows for both open-vocabulary querying and high-quality novel view synthesis in an efficient manner. The method consists of several key components:
3.1.1 3D Gaussian Splatting
The method builds upon the efficient 3D Gaussian Splatting technique [2] for high-quality novel view synthesis, which represents the scene with 3D Gaussians initialized from a sparse point cloud generated via Structure-from-Motion (SfM). Each Gaussian is characterized by its position, anisotropic covariance, opacity, and color via spherical harmonics. Such a representation gives rise to an efficiently differentiable projection to 2D for real-time rendering. The final image derives from the alpha blending model used in volumetric rendering, represented as:
$C=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))c_{i}$ (1)
where $C$ is the accumulated color along the viewing ray, $T_{i}$ is the accumulated transmittance to the $i$-th sample. $\sigma_{i}$ and $\delta_{i}$ denote the density and segment length of the $i$-th Gaussian, and $c_{i}$ is the color contribution of the $i$-th Gaussian.
3.1.2 Dense Language Feature Extraction
Dense language features are extracted from multi-view images using both CLIP [18] and DINO [26] models. These features can provide rich semantic information about the scene.
3.1.3 Feature Quantization
To reduce the high memory requirement of the raw language feature embeddings on dense 3D Gaussians, a quantization scheme [27] is proposed. It creates a discrete feature space, yielding semantic information compression without losing its effectiveness.
3.1.4 Compact Semantic Features on 3D Gaussians
Instead of directly embedding high-dimensional language features, LE3DGS learns compact semantic feature vectors for each 3D Gaussian. These are then decoded into discrete semantic indices using a small MLP.
3.1.5 Adaptive Spatial Smoothing
To handle semantic ambiguities arising from multi-view inconsistencies, LE3DGS introduces an adaptive smoothing mechanism [21]. The smoothing loss is:
$L_{smo}=\|s_{MLP}-s^{*}_{G}\|^{2}+\max(u^{*}_{G},w_{s})\|s^{*}_{MLP}-s_{G}\|^{2}$ (2)
where $s_{MLP}$ are smoothed semantic features generated by an MLP, $s_{G}$ represents the original semantic features of the 3D Gaussians, $u_{G}$ is the learned uncertainty value for each 3D Gaussian, and $w_{s}$ is a minimal weight for the semantic smoothing term. The * operator denotes stop gradient, preventing backpropagation through these terms.
3.1.6 Optimization
It performs a joint optimization over 3D Gaussians for appearance and semantic information. The total loss of the reconstruction and the semantics is:
$\mathcal{L}=\lambda_{s}\mathcal{L}_{s}+\lambda_{smo}\mathcal{L}_{smo}.$ (3)
$\mathcal{L}_{s}$ is the semantic loss:
$\mathcal{L}_{s}=\lambda_{CE}\mathcal{L}_{CE}+\lambda_{u}\mathcal{L}_{u}$ (4)
where $\mathcal{L}_{CE}$ is cross entropy loss and $\mathcal{L}_{u}$ is the uncertainty regularization loss. For further details, the reader is referred to [21].
3.1.7 Inference
During inference, the language feature map $F$ is computed as:
$\mathcal{F}=\mathcal{M}_{\text{infer }}\mathbf{S}$ (5)
where $\mathcal{M}_{\text{infer }}$ is the inferred distribution of semantic indices and $\mathbf{S}$ is the matrix of quantized language features. To detect objects in a 3D scene based on a given text prompt, the process for calculating a relevance map of $\mathcal{F}$ begins with computing the CLIP embedding $\phi_{\texttt{quer}}$ for the text query. Additionally, the CLIP embeddings $\phi^{i}_{\texttt{canon}}$ for a set of predefined canonical phrases is calculated. The final relevancy score for each $\phi_{lang}$ as presented in [17] is given by:
$\text{relevancy}=\min\left(\frac{\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}{\exp(\mathcal{F}\cdot\phi^{i}_{\texttt{canon}})+\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}\right)$ (6)
where $\cdot$ operation is cosine similarity between the two. This scoring mechanism measures how much the rendered embedding is closer to the query embedding with respect to the canonical embedding, indicating the most relevant elements of the scene for the given text prompt. All renderings in [21, 17, 24] use the same predefined canonical phrases: “object”, “things”, “stuff”, and “texture”.
3.2 Inference Using LLM
The canonical phrases serve as a baseline or reference point against which the query and rendered embeddings are compared. They help to contextualize the similarity between the query and the rendered embeddings. Ideally, the canonical phrases should be different from the query and represent a diverse set of concepts or objects that might appear in the scene. They should not be too similar to the query, as that would defeat their purpose.
We can determine if the rendered embedding is more similar to the query than to these baseline concepts by comparing the rendered embedding to both the query and the canonical phrases. If a rendered embedding is significantly nearer to the query than to any canonical phrase, it is probably highly relevant to the query. Conversely, if a rendered embedding is nearer to a canonical phrase than to the query, it likely has little relevance to the query. Accordingly, having a diverse set of canonical phrases helps to cover a wide range of potential scene elements, making the relevancy scoring more robust.
Therefore, we propose leveraging a LLM to generate contextually relevant queries and canonical phrases based on the initial scene description. This approach provides a more adaptive and comprehensive inference framework. We prompt the LLM to identify target objects in the scene, and it responds with:
  • A main query object,
  • Helping positive words: related terms or attributes of the main query, and
  • Canonical phrases: objects to differentiate from the query, including similar objects, nearby objects, and background elements.
This LLM-driven approach allows for dynamic adaptation to diverse driving scenarios, enhancing the system’s ability to understand and interpret complex scenes.
Algorithm 1 quantifies the relevance of language features to a given query, incorporating auxiliary information from helping queries and canonical phrases. It requires $\mathcal{M}$ the language feature map derived from the 3D Gaussians. LLM provides a primary query $p_{\texttt{quer}}$, a set of helping queries $p_{\texttt{help}}$ to provide context, and canonical phrases $p_{\texttt{canon}}$. The process begins by encoding the textual inputs using a CLIP text encoder in Lines 1 to 3. We concatenate query and helping query embeddings to the positive embeddings in Line 4. Then cosine similarities between positive embeddings and the language feature map and between canonical embeddings and the language feature map are computed in Lines 5 and 6. Finally, the relevancy score is calculated using a softmax-like formula in line 7. Notably, the relevancy score represents the probability that each pixel corresponds to the queried object. Subsequently, applying a threshold to these scores allows for the segmentation of the queried elements within the image.
Algorithm 1 Relevancy Score Calculation
0:  language feature map: $\mathcal{M}$, query, helping query, and canonical phrase from LLM: $p_{\texttt{quer}}$, $p_{\texttt{help}}$, $p_{\texttt{canon}}$
0:  relevancy score
1:  $\phi_{\texttt{quer}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{quer}})$
2:  $\phi_{\texttt{help}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{help}})$
3:  $\phi_{\texttt{canon}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{canon}})$
4:  $\phi_{\texttt{pos}}\leftarrow\text{CONCATENATE}(\phi_{\texttt{quer}},\phi_{\texttt{help}})$
5:  $\sigma_{\texttt{pos}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{pos}},\mathcal{M})$
6:  $\sigma_{\texttt{canon}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{canon}},\mathcal{M})$
7:  $\text{score}\leftarrow\min\left(\frac{\exp(\sigma_{\texttt{pos}})}{\exp(\sigma_{\texttt{canon}})+\exp(\sigma_{\texttt{pos}})}\right)$
8:  return  score
4 Experiments
4.1 Dataset
We use a subset of the WayveScenes101 dataset [28], designed specifically to advance novel view synthesis in challenging driving scenarios. We selected a subset of five representative scenes (IDs: 12, 21, 36, 78 and 88) totaling about 500 images. This selection balances computational feasibility with the need for diverse and challenging data, encompassing various environmental conditions, lighting variations, and complex urban environments. This allows us to thoroughly evaluate our approach in handling the complexities of autonomous driving environments. For quantitative evaluation of our method’s semantic understanding capabilities, we generated segmentation maps for the images in the evaluation set using a pre-trained Segformer model [29] on the Cityscapes dataset [30]. This semantic information allows us to assess semantic accuracy.
4.2 Implementation
We trained the LE3DGS using the same loss function as described in [21] on the Wayvesene101 dataset. We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the deafault. The training process involved two main phases: dense semantic feature extraction and 3D Gaussian point optimization. For the semantic feature extraction, we set the number of epochs to 500, ensuring a rich semantic representation of the scene. In the optimization phase for 3D Gaussian points, we employed a 2D position gradient threshold of 0.0001, with a total of 20,000 iterations. The densification process was halted halfway through, at 10,000 iterations. All other hyperparameters were kept consistent with those used in [2]. The entire training process was conducted on a single RTX 4090 GPU, taking approximately 40 minutes to complete for each scene in the dataset. The mean peak signal-to-noise ratio (PSNR) achieved is around 25.14.
Table 1: Input structure for GPT-3.5 Turbo query generation in our implementation. This setup guides the language model to provide structured information about objects relevant to driving scenarios.
Role Content
System A helpful assistant for a computer vision task, providing structured information about objects to pay attention to while driving. Include both important objects and nearby objects that might be at the borders. Instructions for formatting the response, including: Main Positive (object to pay attention to) 1-4 Helping Positives (related terms or attributes) 4-6 Negatives (objects to differentiate from, including similar objects, nearby objects, and background elements)
User "Driving through an intersection in an {Road Type} on a {Weather} {Time of Day}, what objects should the driver pay attention to? " or show the {Object}. "
For the inference phase, we leveraged the API of GPT-3.5 Turbo [31] to dynamically generate queries. For each primary query, the model also provided 1 to 4 helping positive words and 4 to 6 canonical phrases. Table 1 denotes the input structure to the LLM model. The {Road Type}, {Weather}, and {Time of Day} in the user prompts could be provided from the dataset scene metadata [28]. We then calculated the relevancy score using Algorithm 1. To obtain object segmentations, we applied a threshold of 0.5 to the computed relevancy scores. Figure 1 illustrates our inference approach pipeline for scene 88 in Wayvescene101 dataset. The canonical phrases used in the baseline approach are “object”, “things”, “stuff”, and “texture”, consistent with the approach described in [21].
Reconstructed
Predefined $p_{\texttt{canon}}$
Ours w/o $p_{\texttt{help}}$
Ours
Figure 2: Visual comparison of query relevance across different scenes with different approaches. From top to bottom query words are: “trees”, “traffic signs”, “pedestrian”, “sidewalk”, and “cars”.
4.3 Results
Our approach demonstrates significant improvements in open-vocabulary segmentation across various urban driving scenarios. Figure 2 provides a visual comparison, showcasing the progression from overly broad segmentation using predefined canonical phrases to refined results with our full LLM-enhanced method. Our approach accurately identifies and segments various elements such as trees, traffic lights, and sidewalks with high precision, illustrating the benefits of LLM-generated context-aware queries and helping positives.
Inference Method $\mathrm{Accuracy}\uparrow$ $\mathrm{Presicion}\uparrow$ $\mathrm{mIoU}\uparrow$ $\mathrm{mAP}\uparrow$
Predefined Canonical Phrase 0.65 0.33 0.23 0.49
Ours w/o $p_{\texttt{help}}$ 0.84 0.42 0.33 0.53
Ours 0.92 0.58 0.37 0.54
Table 2: Comparison of inference methods: our approach (with and without helping positive words $p_{\texttt{help}}$) versus predefined canonical phrases.
Quantitative results in Table 2 support these visual observations. Our full method achieves the highest performance across all metrics, with notable improvements in Accuracy (0.92), Precision (0.58), and mean Intersection over Union (mIoU) (0.37). These represent significant gains over the predefined canonical phrase method (0.65, 0.33, and 0.23 respectively). The ablation study (Ours w/o $p_{\texttt{help}}$) indicates the robustness of our LLM-driven query generation, outperforming the baseline even without helping positive words.
(a) Mean Accuracy
(b) Mean IoU
(c) Mean Precision
Figure 3: Performance comparison of our methods against baseline across different object classes for the scenes in dataset.
Figure 3 further breaks down performance across various urban scene elements. Our method shows significant improvements in Mean IoU and Mean Precision for challenging classes like "sidewalk", "traffic light", and "traffic signs", while maintaining an edge in common classes such as "building" and "car". These results collectively demonstrate the effectiveness of our LLM-enhanced approach in improving scene understanding and object segmentation across diverse urban environments.
For more detailed qualitative results, please refer to section A.1 in the appendix.
5 Discussion
Our approach demonstrates significant improvements in open-vocabulary 3D scene understanding, but there are several important aspects to consider:
Limitations of 3D Gaussian Splatting
While 3D Gaussian Splatting is effective for certain scenes, it struggles with generating large and dynamic scenes [32]. This limitation can affect the overall performance and applicability of our approach in more dynamic environments, such as busy urban intersections or highways with fast-moving vehicles.
System Integration and Real-time Performance
The integration of our approach with existing autonomous driving systems presents both opportunities and challenges. While our method offers enhanced semantic understanding, its real-time performance in the context of rapidly changing driving scenarios requires further optimization. Current autonomous systems often rely on faster, albeit less semantically rich, perception methods. Bridging this gap involves not only improving our method’s computational efficiency but also developing strategies for seamless integration with existing sensor fusion and decision-making pipelines.
6 Conclusion
In this study, we proposed a novel method that integrates LLMs with LE3DGS to improve autonomous vehicle scene understanding. By leveraging the robust reasoning capabilities of LLMs and the detailed spatial representation of 3D Gaussians, our approach significantly improves the accuracy and precision of open vocabulary object segmentation and scene interpretation tasks. Our experimental results indicate that our method outperforms traditional predefined canonical phrase-based approaches on all aspects of accuracy, precision, and mean Intersection over Union.
Future work will focus on enhancing our system’s capability to handle dynamic scenes by integrating LLMs with advanced 3D Gaussian Splatting techniques. We aim to fuse our approach with multi-modal sensor data, such as LiDAR, to create a more robust and comprehensive perception system. Another crucial next step is to seamlessly integrate our advanced semantic understanding with autonomous driving planners. This integration will involve developing efficient interfaces that translate our rich semantic insights into actionable information, directly informing path planning and decision-making algorithms for safer and more intelligent autonomous navigation.
Appendix A Appendix
A.1 Qualitative Results
Figures 4, 5, and 6 showcase our approach’s effectiveness in open vocabulary segmentation across diverse driving scenes. Each figure presents, from top to bottom: ground truth images, Segformer-generated segmentation maps (used for quantitative evaluation), our high-quality scene reconstructions using LE3DGS, and our open vocabulary segmentation results. Our method accurately highlights key elements like cars, pedestrians, trees, buildings, and traffic signs/lights. For quantitative evaluation, we align with Segformer classes [33], but our visualizations include additional classes to demonstrate the flexibility of our approach. These figures illustrate our method’s capability to accurately reconstruct and semantically segment complex urban scenes, even for classes beyond traditional segmentation models.
Figure 4: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scene 88.
(a) Scene 12
(b) Scene 21
Figure 5: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 12 and 21.
(a) Scene 36
(b) Scene 78
Figure 6: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 36 and 78.